Can you help me with this word problem?
The function f(x) = 2x2 +22x + 320 models the number of inmates in federal and state prisons in the United States [ f (x), in thousands], x years after 1980. Find and interpret f(14).
To start is this right?
4 + 22x + 320 = 22x + 324
-22 -22
x = 312

Answer by checkley71(8403)

I BELIEVE YOU ARE MISINTERPRETING THE X FACTOR; X MUST BE REPLACED BY 14.
THUS THE 2X2 TERM SHOULD BE INTERPRETED AS (2*14^2) NOT (2*2).
ALSO YOU CANNOT SUBTRACT 22 FROM THE TERMS 22X & 324.
ALTHOUGH YOU CAN DIVIDE OR MULTIPLY THESE TERMS BY 22, IF IT IS APPROPRIATE.
f(x)=2x^2+22x+320
f(14)=2*14^2+22*14+320
f(14)=2*196+308+320
f(14)=392+308+320
f(14)=1,020